# Janus Skyrmion: Interfacial Quasiparticle with Two-Faced Helicity

### v1

@@ Abstract
Current-induced dynamics.—The Janus skyrmion is an interfacial state stabilized at the interface. Although it is a two-dimensional topological spin texture, its dynamics will be affected by the interface and should be largely different to that of a normal skyrmion. Here, we explore the dynamics of a Janus skyrmion at the interface driven by a vertical spin current [13, 14, 15]. Initially, a Janus skyrmion is relaxed at the center of a magnetic layer ($100\times 100\times 1$ nm3) with periodic boundary conditions applied in both $x$ and $y$ directions. The DM interaction in the left half of the magnetic layer is of Néel-type, and it is of Bloch-type in the right half. We then apply a vertical spin current to drive the Janus skyrmion into motion, which can be realized by the spin Hall effect in the substrate underneath the magnetic layer [13, 14, 15, 16, 17].
@@
In stark contrast to a normal skyrmion driven by the spin Hall torque, of which the velocity direction (i.e., skyrmion Hall angle) depends on $\theta$ but the speed is independent of $\theta$, the current-induced dynamics of the Janus skyrmion is limited to a one-dimensional motion along the interface (i.e., $\pm y$ directions) without showing the skyrmion Hall effect. Both the speed and motion direction of the Janus depend on $\theta$, as shown in Figs. 3(a) and 3(b) (Supplementary Videos 1 and 2 [48]). It can move toward the +y+y or y-y direction, and its speed varies with $\theta$ for a given $j$.
@@
A normal skyrmion is usually repelled by magnetic surfaces and interfaces, while the Janus skyrmion is an interfacial state attracted by its hosting interface. Hence, the Janus skyrmion can be viewed as a unique topological quasiparticle strictly confined in a one-dimensional potential well [50]. Its current-induced dynamics could be described by the Thiele equation [51, 27, 17, 16] with the assumption that an interface-induced force $\boldsymbol{F}_{x}$ is acting on the Janus skyrmion and $v_{x}=0$. The velocity solution is thus given as $v_{y}=c\cdot\cos{(\theta+\pi/4)}$ with $c=\left|\sqrt{2}u/2\alpha{\mathcal{D}}\right|\cdot{\mathcal{I}}$ being a constant determined by the driving force $u\sim j$ and the Janus skyrmion profile (Supplementary Note 1 [48]). Note that $\mathcal{D}$} is the diagonal entries of the dissipative tensor in the Thiele equation and $\mathcal{I}$} is related to the spin-torque efficiency over the skyrmion [51, 27, 17, 16]. In Fig. 3, it can be seen that the simulation results are in a good agreement with the theoretical velocity solutions, where $c=20$ for $j=1$ MA cm-2 and $c=60$ for $j=3$ MA cm-2. Note that a current-driven Janus skyrmion may be detached from the interface when $\boldsymbol{F}_{x}$ is too strong (Supplementary Fig. 2 [48]), which then transforms into a normal skyrmion (Supplementary Videos 4 and 5 [48]). We also compare the current-induced motion of a Janus skyrmion with that of normal Néel-type and Bloch-type skyrmions in Supplementary Note 1 [48].
@@
One-dimensional random walk.—As the Janus skyrmion can only move along the interface, it should be able to demonstrate one-dimensional random walk at the interface. To examine this prediction, we introduce a thermal fluctuation term into the LLG equation [52], and simulate the Brownian motion of a Janus skyrmion at finite temperature. The spin dynamics at finite temperature is governed by the stochastic LLG equation $\partial_{t}\boldsymbol{m}=-\gamma_{0}\boldsymbol{m}\times(\boldsymbol{h}_{\text{eff}}+\boldsymbol{h}_{\text{f}})+\alpha(\boldsymbol{m}\times\partial_{t}\boldsymbol{m})$, where $\boldsymbol{h}_{\text{f}}$ is a thermal fluctuating field satisfying $h_{i}(\boldsymbol{x},t)>=0$ and $h_{i}(\boldsymbol{x},t)h_{k}(\boldsymbol{x}^{\prime},t^{\prime})>=\frac{2\alpha k_{\text{B}}T}{M_{\text{S}}\gamma_{0}\mu_{0}V}\delta_{ik}\delta(\boldsymbol{x}-\boldsymbol{x}^{\prime})\delta(t-t^{\prime})$. $i$ and $k$ are Cartesian components, $k_{\text{B}}$ is the Boltzmann constant, $T$ is the temperature, and $V$ is the volume of a single mesh cell. $\delta_{ik}$ and $\delta(\dots)$ denote the Kronecker and Dirac delta symbols, respectively. In all simulations, a fixed finite-temperature integration time step of $10$ fs is applied.
@@
In Fig. 5, it can be seen that the Janus skyrmion demonstrates obvious one-dimensional random walk along the interface (Supplementary Videos 6-8 [48]). The random walk diffusion in the $y$ dimension increases with $T$ [Fig. 5(b)], but the skyrmion position in the $x$ dimension only slightly fluctuates with time [Fig. 5(a)]. The Janus skyrmion size also increases with $T$, which can be seen from time-dependent $m_{z}$ in Fig. 5(c). The one-dimensional Brownian motion is a unique thermal dynamic feature of interfacial quasiparticles. A normal skyrmion usually demonstrates two-dimensional Brownian motion [53, 54, 55, 56, 57, 58, 59, 60, 61].
@@
Conclusions.—We have explored the unique dynamics of a Janus skyrmion induced by magnetic field, spin current, or temperature. The Janus skyrmion is stabilized at the interface between two magnetic regions with different types of antisymmetric exchange interactions. Its helicity structure include both Néel-type and Bloch-type components. The dynamic behaviors of a Janus skyrmion largely differ from that of a normal skyrmion as it is topological quasiparticle stabilized and attached to the interface. The dynamic degree of freedom of a Janus skyrmion is limited by the interface, and consequently, the speed of a Janus skyrmion moving along the interface could be controlled by the spin polarization angle, and it will not show the skyrmion Hall effect. Thermal fluctuations also lead to one-dimensional random walk of the Janus skyrmion at the interface, which is different to normal skyrmions that can diffuse in two dimensions. It is also found that both in-plane and out-of-plane magnetic fields can modify the size of a Janus skyrmion due to its helicity asymmetry. The size of normal skyrmions with symmetric helicity structures cannot be adjusted by an in-plane magnetic field. The asymmetric helicity structures of heart-shaped Janus skyrmions result in distinctive responses to magnetic field, spin current, and temperature. As the Janus skyrmion does not show the skyrmion Hall effect, it could be a promising building block for racetrack-type device applications [62, 63, 64, 65, 66]. This work will pave the way for the understanding of topological quasiparticles at material interfaces.
@@ Acknowledgements.
X.Z. and M.M. acknowledge support by the CREST, the Japan Science and Technology Agency (Grant No. JPMJCR20T1). X.Z. also acknowledges support by the Grants-in-Aid for Scientific Research from JSPS KAKENHI (Grant No. JP25K17939 and No. JP20F20363). M.M. also acknowledges support by the Grants-in-Aid for Scientific Research from JSPS KAKENHI (Grants No. JP25H00611, No. JP24H02231, No. JP23H04522, and No. JP20H00337) and the Waseda University Grant for Special Research Projects (Grant No. 2025C-133). R.Z. acknowledges support by the State Key Laboratory of Displays and Opto-Electronics (Project Reference: ITC-PSKL12EG02). C.R. and C.J.O.R. acknowledge the support by the U.S. Department of Energy through the Los Alamos National Laboratory. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of the U.S. Department of Energy (Contract No. 892333218NCA000001). Y.Z. acknowledges support by the Shenzhen Fundamental Research Fund (Grant No. JCYJ20210324120213037), the National Natural Science Foundation of China (12374123), the Guangdong Basic Research Center of Excellence for Aggregate Science, and the 2023 SZSTI Stable Support Scheme.

### v2

@@ II.3 Current-induced dynamics
The Janus skyrmion is an interfacial state stabilized at the interface. Although it is a 2D topological spin texture, its dynamics will be affected by the interface and should be largely different to that of a conventional skyrmion. Here, we explore the dynamics of a Janus skyrmion at the interface driven by a vertical spin current [16, 17, 18]. Initially, a Janus skyrmion is relaxed at the center of a magnetic layer ($100\times 100\times 1$ nm3) with periodic boundary conditions applied in both $x$ and $y$ directions. The DM interaction in the left half of the magnetic layer is of Néel-type, and it is of Bloch-type in the right half. We then apply a vertical spin current to drive the Janus skyrmion into motion, which can be realized by the spin Hall effect in the substrate underneath the magnetic layer [16, 17, 18, 19, 20].
@@
In stark contrast to a conventional skyrmion driven by the spin Hall torque, of which the velocity direction (i.e., the skyrmion Hall angle) depends on $\theta$ but the speed is independent of $\theta$, the current-induced dynamics of the Janus skyrmion is limited to a 1D motion along the interface (i.e., $\pm y$ directions) without showing the skyrmion Hall effect [14, 17, 18]. Both the speed and motion direction of the Janus depend on $\theta$, as shown in Figs. 3(a) and 3(b) (Supplementary Videos 1 and 2 [66]). It can move toward the +y+y or y-y direction, and its speed varies with $\theta$ for a given $j$.
@@
A conventional skyrmion is usually repelled by magnetic surfaces and interfaces, while the Janus skyrmion is an interfacial state attracted by its hosting interface. Hence, the Janus skyrmion can be viewed as a unique topological quasiparticle strictly confined in a 1D potential well [68]. Its current-induced dynamics could be described by the Thiele equation [69, 40, 20, 19] with the assumption that an interface-induced force $\boldsymbol{F}_{x}$ is acting on the Janus skyrmion and $v_{x}=0$. The velocity solution is thus given as $v_{y}=c\cdot\cos{(\theta+\pi/4)}$ with $c=\left|\sqrt{2}u/2\alpha{\mathcal{D}}\right|\cdot{\mathcal{I}}$ being a constant determined by the driving force $u\sim j$ and the Janus skyrmion profile (Supplementary Note 1 [66]). Note that $\mathcal{D}$} is the diagonal entries of the dissipative tensor in the Thiele equation and $\mathcal{I}$} is related to the spin-torque efficiency over the skyrmion [69, 40, 20, 19]. In Fig. 3, it can be seen that the simulation results are in a good agreement with the theoretical velocity solutions, where $c=20$ for $j=1$ MA cm-2 and $c=60$ for $j=3$ MA cm-2. Note that a current-driven Janus skyrmion may detach from the interface when $\boldsymbol{F}_{x}$ is too strong (Supplementary Fig. 3 [66]), which then transforms into a conventional skyrmion (Supplementary Videos 4 and 5 [66]). Besides, the imperfections at the interface may reduce the Janus skyrmion speed and induce instability, especially at high current densities (Supplementary Fig. 4 [66]). We also compare the current-induced motion of a Janus skyrmion with that of conventional Néel-type and Bloch-type skyrmions in Supplementary Note 1 [66].
Figure 5: 1D Brownian random walk of a Janus skyrmion at the interface. (a) Time-dependent position of the Janus skyrmion in the $x$ dimension. (b) Time-dependent position of the Janus skyrmion in the $y$ dimension. (c) Time-dependent $m_{z}$ of the model. Here, the Brownian dynamics of the Janus skyrmion is simulated at three different temperatures ($T=10$, $50$, and $100$ K). See Fig. 3 caption for other parameters.
II.4 One-dimensional random walk
@@
As the Janus skyrmion can only move along the interface, it should be able to demonstrate 1D random walk at the interface. To examine this prediction, we introduce a thermal fluctuation term into the LLG equation [70], and simulate the Brownian motion of a Janus skyrmion at finite temperature. The spin dynamics at finite temperature is governed by the stochastic LLG equation $\partial_{t}\boldsymbol{m}=-\gamma_{0}\boldsymbol{m}\times(\boldsymbol{h}_{\text{eff}}+\boldsymbol{h}_{\text{f}})+\alpha(\boldsymbol{m}\times\partial_{t}\boldsymbol{m})$, where $\boldsymbol{h}_{\text{f}}$ is a thermal fluctuating field satisfying $h_{i}(\boldsymbol{x},t)>=0$ and $h_{i}(\boldsymbol{x},t)h_{k}(\boldsymbol{x}^{\prime},t^{\prime})>=\frac{2\alpha k_{\text{B}}T}{M_{\text{S}}\gamma_{0}\mu_{0}V}\delta_{ik}\delta(\boldsymbol{x}-\boldsymbol{x}^{\prime})\delta(t-t^{\prime})$. $i$ and $k$ are Cartesian components, $k_{\text{B}}$ is the Boltzmann constant, $T$ is the temperature, and $V$ is the volume of a single mesh cell. $\delta_{ik}$ and $\delta(\dots)$ denote the Kronecker and Dirac delta symbols, respectively. In all simulations, a fixed finite-temperature integration time step of $10$ fs is applied.
@@
In Fig. 5, it can be seen that the Janus skyrmion demonstrates obvious 1D random walk along the interface (Supplementary Videos 6-8 [66]). The random walk diffusion in the $y$ dimension increases with $T$ [Fig. 5(b)], but the skyrmion position in the $x$ dimension only slightly fluctuates with time [Fig. 5(a)]. The Janus skyrmion size also increases with $T$, which can be seen from time-dependent $m_{z}$ in Fig. 5(c). The 1D Brownian motion is a unique thermal dynamic feature of interfacial quasiparticles. A conventional skyrmion usually demonstrates 2D Brownian motion [71, 72, 73, 74, 75, 76, 77, 78, 79].
III Conclusion
@@
We have explored the unique dynamics of a Janus skyrmion induced by magnetic field, spin current, or temperature. The Janus skyrmion is stabilized at the interface between two magnetic regions with different types of antisymmetric exchange interactions. Its helicity structure include both Néel-type and Bloch-type components. The dynamic behaviors of a Janus skyrmion largely differ from that of a conventional skyrmion as it is topological quasiparticle stabilized and attached to the interface. The dynamic degree of freedom of a Janus skyrmion is limited by the interface, and consequently, the speed of a Janus skyrmion moving along the interface could be controlled by the spin polarization angle, and it will not show the skyrmion Hall effect. The skyrmion Hall effect is a dynamic feature intrinsic to conventional ferromagnetic skyrmions [13, 14, 17, 18, 19, 20, 21, 22, 42], which could be avoided in antiferromagnetic systems [62, 63, 80, 71, 81]. Moreover, thermal fluctuations lead to 1D random walk of the Janus skyrmion at the interface, which is different to conventional skyrmions that can diffuse in two dimensions. It is also found that both in-plane and out-of-plane magnetic fields can modify the size of a Janus skyrmion due to its helicity asymmetry. The size of conventional skyrmions with symmetric helicity structures cannot be adjusted by an in-plane magnetic field. The asymmetric helicity structures of heart-shaped Janus skyrmions result in distinctive responses to magnetic field, spin current, and temperature. As the Janus skyrmion does not show the skyrmion Hall effect, it could be a promising building block for racetrack-type device applications [82, 83, 84, 85, 86]. This work will pave the way for the understanding of interfacial topological quasiparticles.
@@ Acknowledgements.
X.Z. and M.M. acknowledge the support by the CREST, the Japan Science and Technology Agency (Grant No. JPMJCR20T1). X.Z. acknowledges the support by the Grants-in-Aid for Scientific Research from JSPS KAKENHI (Grants No. JP25K17939 and No. JP20F20363). R.Z. acknowledges the support by the State Key Laboratory of Displays and Opto-Electronics (Project Reference: ITC-PSKL12EG02). Q.S. acknowledges the support by the RGC General Research Fund (No. 16309924). Y.Z. acknowledges the support by the Shenzhen Fundamental Research Fund (Grant No. JCYJ20210324120213037), the National Natural Science Foundation of China (Grant No. 12374123), the Guangdong Basic Research Center of Excellence for Aggregate Science, and the 2023 SZSTI Stable Support Scheme. C.R. and C.J.O.R. acknowledge the support by the U.S. Department of Energy through the Los Alamos National Laboratory. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of the U.S. Department of Energy (Contract No. 892333218NCA000001). M.M. acknowledges the support by the Grants-in-Aid for Scientific Research from JSPS KAKENHI (Grants No. JP25H00611, No. JP24H02231, No. JP23H04522, and No. JP20H00337) and the Waseda University Grant for Special Research Projects (Grant No. 2025C-133).